# Sub-shot noise sensitivity via deformed four-headed kitten states

### v2

@@ III.2 Fluctuations in photon numbers
The photon addition and subtraction operations, as examined in our case, are anticipated to affect the photon number distribution (PND) of our quantum states. For example, in the compass state, the sizes of sub-Planck structures, phase-space sensitivity, and PND are all proportional to the macroscopic parameter. We now illustrate the effects of multi-photon operations on the photon statistics of resulting deformed coherent states To investigate the PND in these SA and AS scenarios of the coherent state, we use the mathematical expression for PND of a quantum state |$\ket{\psi}$, denoted as $P_{\ket{\psi}}(n):=\left|\braket{n}{\psi}\right|^{2}$, where |$\ket{n}$ represents the Fock state [7].
The photon number distributions for the SA and AS cases, denoted as $P_{\ket{\oplus}}(n)$ and $P_{\ket{\ominus}}(n)$, are evaluated as follows:
$P_{\ket{\oplus}}(n)=\frac{N_{\oplus}[(q+n)!]^{2}\kappa}{n![(q+n-r)!]^{2}}~{}\text{and}~{}P_{\ket{\ominus}}(n)=\frac{N_{\ominus}n!\kappa}{[(n-r)!]^{2}}$ (19)
@@
Let us now analyze these distributions. Here, Figs. 6(a)-(c) show the PND for the SA case of the coherent state for different situations based on the varying amount of added and subtracted photons. As depicted in Fig. 6(a), PND with $r=0$ and $q=0$ exhibits a Poissonian distribution, which obviously corresponds to a coherent state. In Fig. 6(b), when $r=4$ photons are added and set $q=0$, the Poissonian distribution shifts to higher values of $n$, with the peak now occurring at a larger mean photon number. In Fig. 6(c), the Poissonian distribution shifts to lower values of $n$, as with an increment in the subtracting photons to $q=4$, and the number of added photons is kept at $r=4$ as in the prior instance. These cases reflects that in the SA examples, a higher $r$ corresponds to a higher average photon number, whereas a higher $q$ corresponds to a lower average photon number in the resultant state.
Figure 7: The overlap function $|F_{\square}(c_{0})|^{2}$ is represented as a blue dashed curve and $|F_{\blacksquare}(c_{0})|^{2}$ as a red solid line, with $r=4$ and $q=4$. The horizontal black dashed line depicts the case where $r=0$ and $q=0$; consequently, the overlap is simply between two compass states.
The PND for AS case is shown in Figs. 6(d)-(f) with a few $r$ and $q$ values. As illustrated in Fig. 6(d), for $q=2$ and $r=0$, the statistics of PND stay constant, showing that raising the number of subtracted photons has no influence on the average photon number of the states when applied directly to a coherent state. As observed in Fig. 6(e), increasing the number of added photons while maintaining the number of subtracted photons constant, that is, with $r=4$ and $q=2$, drives the Poissonian distribution to a larger $n$. In Fig. 6(f), the number of subtracted photons is increased to $q=4$ without changing $r$, and it is observed that subtracting photons from a coherent state has no influence on the related PND. This is an additional confirmation of how photon subtraction operations keeps PND of this case invariant. Hence, photon addition in both the SA and AS situations increases the average photon number in the subsequent states; however, the AS case maintains the distribution at higher mean photon values, implying that this situation has higher average photon numbers than the SA case.
Figure 8: (a)-(c) show the Wigner function of the SA case of the four-headed kitten state, and (d)-(f) show the analogous AS cases. (a) $r=12$, $q=12$, (b) $r=24$, $q=12$, (c) $r=24$, $q=20$, (d) $r=12$, $q=12$, (e) $r=24$, $q=12$, and (f) $r=24$, $q=20$. In all cases, $c_{0}=1$.
@@ III.3 Impact of multiphoton operations
In §III.1, we noticed that the SA and AS variants of a standard coherent state exhibit fascinating phase-space properties that surpass those of the original coherent states. It is observed that the sequence in which photon operations are applied significantly affects the outcome, resulting in two distinct quantum states with different phase-space characteristics. We now extended this basic notion to the main quantum states of the present work.
First, for our SA scenario, $r$ photons are added to the kitten state $\ket{\diamondsuit}$ as given in Eq. (5), followed by the subtraction of $q$ photons. In the AS scenario, $q$ photons are first subtracted from the compass state, and then $r$ photons are added. These photon operations on the kitten state produce new quantum states with distinct phase-space characteristics.
The resulting SA case is denoted as $\ket{\square}$ and mathematically described as:
@@
denotes the normalization coefficient.
In the same way, the AS situation, denoted as $\ket{\blacksquare}$, is represented as follows:
@@
$N_{\blacksquare}=$ $\sum^{4}_{i,j=1}\sum^{r}_{n=0}(-1)^{n}\Gamma G_{\alpha_{i},\alpha_{j}}\text{e}^{\alpha^{*}_{i}\alpha_{j}}(\alpha_{i}\alpha_{j})^{q}$
$(\alpha_{j})^{r-n}(\alpha^{*}_{i})^{r-n}$ (24)
represents the normalization factor for this case. When photon subtraction is combined with a photon addition operator, the effect of the parameter $q$ is highly apparent, as we will show later in the following discussion.
We now compare the original compass state with the proposed variants by assessing their overlap. This overlap indicates how distinct the proposed states are from the original compass state. The overlap between our SA case and the compass state is represented by $F_{\square}(c_{0}):=\braket{\square}{\diamondsuit}$ and is calculated as
$F_{\square}(c_{0})=(-\text{i})^{r+q}\left[N_{\square}N_{\diamondsuit}\right]^{\nicefrac{{-1}}{{2}}}\sum^{4}_{i,j=1}F_{\ket{\square_{i}}\bra{\square_{j}}}$ (25)
@@
$F_{\ket{\square_{i}}\bra{\square_{j}}}:=G_{\alpha_{i},\alpha_{j}}\text{e}^{\alpha^{*}_{i}\alpha_{j}}H_{r,q}\left[\text{i}\alpha^{*}_{i},\text{i}\alpha_{j}\right],$ (26)
and is depicted with a blue dashed line in Fig. 7.
Now, we provide the overlap between AS and compass state, which is denoted as $F_{\blacksquare}(c_{0}):=\braket{\blacksquare}{\diamondsuit}$. This overlap reads
$F_{\blacksquare}(c_{0})=\left[N_{\blacksquare}N_{\diamondsuit}\right]^{\nicefrac{{-1}}{{2}}}\sum^{4}_{i,j=1}F_{\ket{\blacksquare_{i}}\bra{\blacksquare_{j}}}$ (27)
with
$F_{\ket{\blacksquare_{i}}\bra{\blacksquare_{j}}}:=G_{\alpha_{i},\alpha_{j}}\text{e}^{\alpha^{*}_{i}\alpha_{j}}(\alpha^{*}_{i})^{q}(\alpha_{j})^{r},$ (28)
and is plotted with a red solid line in Fig. 7.
In Fig. 7, the black dotted line represents the overlap between two same compass states, and recalling that the parameter $c_{0}$ measures the distance from the origin of the phase space, as discussed in §II.2, this parameter has a strong impact on the characteristics of the original state, that is, the existence of the sub-Planck structures, and enhanced sensitivity is associated with higher $c_{0}$ values. For example, the scenario represented in Fig. 1(a) for comparatively smaller values of $c_{0}$ does not have the capacity to exhibit these traits, and this particular case is named as a four-headed kitten state in the present work.
Note that in our plots, we only investigate the overlaps for the certain situation when $r=q=4$, because it is understood here that for other higher values of $r$ and $q$, our proposed states keep the difference than the compass state provided that $c_{0}$ is smaller, as depicted here; simply, this result is meaningful for other settings of $r$ and $q$. In these plots, it is clear that greater $c_{0}$ values push the blue dotted and red solid lines closer to the horizontal black dotted line, implying that the SA and AS cases revived to the original compass state. In our work, we set $c_{0}=1$ and investigate the corresponding phase-space characteristics of the four-headed cat state using multiple options of photon addition and subtraction operations. Here, we simply refer to these consequent states as SA and AS cases of this kitten state. Here, note that our AS case is the eigenstate of the operator $\hat{a}^{4}$, i.e., $\hat{a}^{4}\ket{\blacksquare}=\ket{\diamondsuit}$. This shows that choosing the parameter $q$ as a multiple of 4 ($q=4k$ with $k\in\mathbb{N}^{+}$) corresponds to a photon-added version of the compass state. When $p=0$, this reduces to the compass state $\ket{\diamondsuit}$ itself.
Figure 9: Zoom in on the central phase-space structures of the cases shown in Fig. 8. The cases on the left are SA, whereas the cases on the right are AS. (a) $r=12$, $q=12$; (b) $r=24$, $q=12$; (c) $r=24$, $q=20$; (d) $r=12$, $q=12$; (e) $r=24$, $q=12$ (f) $r=24$, $q=20$. In each example, $c_{0}=1$.
Figure 10: The zoom in on the center phase-space structures of the AS case of the four-headed kitten state with multiple parameter $r$ and $q$ values. (a) $r=1$, $q=1$; (b) $r=5$, $q=5$; (c) $r=9$, $q=9$; (d) $r=16$, $q=10$; (e) $r=22$, $q=10$; (f) $r=22$, $q=18$. For all cases, $c_{0}=1$.
@@ III.4 Unique phase-space features
In the previous section, we have introduced the SA and AS cases of our interest, each producing distinct quantum states. In these instances, photon addition and subtraction processes appeared to have a main impact on their phase-space characteristics, and in this section, we particularly employ the Wigner function to investigate the corresponding phase space of these quantum states. To obtain the Wigner function for each case, Eq. (1) is employed, and we denote $W_{\ket{\square}}(\beta)$ and $W_{\ket{\blacksquare}}(\beta)$ as the corresponding Wigner functions of SA and AS cases of the kitten state, respectively.
Let us now examine the Wigner distributions for each scenario. For the SA case of our four-headed kitten state introduced in Eq. (21), the Wigner distribution is calculated as follows:
@@
×$\times(\alpha_{j}-2\beta^{*})^{r-n}.$ (33)
The relevant Wigner functions are presented in Figs. 8, with Fig. 9 illustrating the zoom in of the central phase-space features of each case, where for both figures the cases labelled with (a)-(c) exhibit the SA, while (d)-(f) provides equivalent AS cases of the kitten state. With specific selections of the parameters $r$ and $q$, Fig. 10 offers a further illustration of the Wigner function of corresponding AS scenarios.
@@
It is readily apparent that our SA and AS instances achieve substantially distinct phase-space characteristics as compared to the four-headed kitten state. Interestingly, a central sub-Planck structure is identified in each of the cases outlined. We focus on the significance of the sub-Planck structure in these states and examine this particular phase-space feature in detail.
@@
Photon addition and subtraction appeared to have significant effects on the Wigner distribution of the corresponding states. Note that the original four-headed kitten state, represented in Fig. 1(a) with $c_{0}=1$, does not exhibit a sub-Planck structure. This implies that the appearance of the sub-Planck structures in our cases is attributed to the photon operations involved. Specifically, it is observed that as the parameter $r$, which represents the number of added photons, increases, the size of the sub-Planck structure reduces uniformly in both the SA and AS cases. This is evident by comparing the scenarios presented in Figs. 9(a) and 9(b) for the SA case, and then Figs. 9(d) and 9(e) are for corresponding AS cases, where increasing $r$ with constant $q$, clearly depicts this impact. Furthermore, increasing the number of photon subtractions $q$ increases the size of the sub-Planck structure in the SA case, as observed by comparing cases depicted in Figs. 8(b) and 8(c), while it has no effect in the AS case, as shown in Figs. 8(e) and 8(f). Note that all the examples of AS described in Figs. 8 and 9 correspond to the merely photon-added case of the compass state, as for these situations $q$ are the multiple of 4, where the corresponding four-headed kitten state is the eigenstate of the operator $\hat{a}^{4}$.
@@
We find that photon addition generally reduces the size of sub-Planck structures in both SA and AS scenarios, while photon subtraction increases their size in the SA case but has no effect on their size in the AS case. To further illustrate this for the AS case, we examine situations where the number of photon subtractions is not a multiple of 4, as shown in Fig. 10. Increasing the photon addition to the state as previously noted also reduces the size of the sub-Planck structures for this specific illustration (Figs. 10(d) and 10(e)), while as expected, varying the photon subtraction in these cases has no impact on the size of the sub-Planck structures (Figs. 10(e) and 10(f)), but in some specific instances, photon subtraction slightly enhances the isotropy of the sub-Planck structures. For example, an enhancement in the isotropy of the central sub-Planck structures of these cases is clearly evident in Figs. 10(a) to 10(c), as demonstrated by the circular structure around the origin. This indicates that, for these scenarios, the sub-Planck structure is uniformly constrained in all directions of the phase space.
Figure 11: Overlap between a state and its slightly translated version with (a)-(c) depict SA examples, whereas (d)-(i) exhibit comparable AS cases: (a) $r=12$, $q=12$, (b) $r=24$, $q=12$, (c) $r=24$, $q=20$, (d) $r=12$, $q=12$, (e) $r=24$, $q=12$, (f) $r=24$, $q=20$, (g) $r=q=1$, (h) $r=q=5$, and (i) $r=q=9$. In all situations, we set $c_{0}=1$ and normalize the intensity plots to unity.

### v3

@@ III.2 Fluctuations in photon numbers
In this section, we explore how the order in which photon addition and subtraction operations are applied to a state influences the photon number statistics. The photon addition and subtraction operations, as examined in our case, are anticipated to affect the photon number distribution (PND) of our quantum states. For example, in the compass state, the sizes of sub-Planck structures, phase-space sensitivity, and the PND are all proportional to the macroscopic parameter $c_{0}$ [47]. We now demonstrate the consequences of photon operations with varying values $r$ and $q$, and for this we evaluate the corresponding PND of each case of reduced coherent states. To investigate the PND in these SA and AS scenarios of the coherent state, we use the mathematical expression for the PND of a quantum state |$\ket{\psi}$, defined as $P_{\ket{\psi}}(n):=\left|\braket{n}{\psi}\right|^{2}$, where |$\ket{n}$ represents the Fock state [7].
The PND for the SA and AS cases, denoted as $P_{\ket{\oplus}}(n)$ and $P_{\ket{\ominus}}(n)$, are evaluated as
$P_{\ket{\oplus}}(n)=\frac{N_{\oplus}[(q+n)!]^{2}\kappa}{n![(q+n-r)!]^{2}}$ (21)
and
$P_{\ket{\ominus}}(n)=\frac{N_{\ominus}n!\kappa}{[(n-r)!]^{2}}$ (22)
@@
Let us now analyze these distributions. Figures 6(a)-6(c) show the PNDs for the SA case of the coherent state for different situations based on the varying amount of added and subtracted photons. As depicted in Fig. 6(a), the PND with $r=0$ and $q=0$ exhibits a Poissonian distribution, which obviously corresponds to a coherent state. In Fig. 6(b), when $r=4$ photons are added and we set $q=0$, the Poissonian distribution shifts to higher values of $n$, with the peak now occurring at a larger mean photon number. In Fig. 6(c), the Poissonian distribution shifts to lower values of $n$, with an incremental subtraction of photons to $q=4$, and the number of added photons is kept at $r=4$ as in the prior instance. These cases reflects that in the SA examples, a higher $r$ corresponds to a higher average photon number, whereas a higher $q$ corresponds to a lower average photon number in the resultant state.
Figure 8: [(a)-(c)] the Wigner function of the SA case of the four-headed kitten state and [(d)-(f)] the analogous AS cases for (a) $r=12$ and $q=12$, (b) $r=24$ and $q=12$, (c) $r=24$ and $q=20$, (d) $r=12$ and $q=12$, (e) $r=24$ and $q=12$, and (f) $r=24$ and $q=20$. In all cases, $c_{0}=1$.
Figure 9: Close-up of the central phase-space structures of the cases shown in Fig. 8. [(a)-(c)] SA and [(d)-(f)] AS for (a) $r=12$ and $q=12$, (b) $r=24$ and $q=12$; (c) $r=24$ and $q=20$, (d) $r=12$ and $q=12$, (e) $r=24$ and $q=12$ (f) $r=24$ and $q=20$. In each example, $c_{0}=1$.
The PND for the AS case is shown in Figs. 6(d)-6(f) for a few $r$ and $q$ values. As illustrated in Fig. 6(d), for $q=2$ and $r=0$, the statistics of the PND stay invariant, showing that raising the number of subtracted photons has no influence on the average photon number of the states when applied directly to a coherent state. As observed in Fig. 6(e), increasing the number of added photons while maintaining the number of subtracted photons constant, that is, with $r=4$ and $q=2$, drives the Poissonian distribution to a larger $n$. In Fig. 6(f), the number of subtracted photons is increased to $q=4$ without changing $r$, and it is observed that subtracting photons from a coherent state has no influence on the related PND. This is an additional confirmation of how photon subtraction operations keep the PND of this case invariant. Hence, photon addition in both the SA and AS situations increases the average photon number in the coherent states; however, the AS case maintains the distribution at higher mean photon values, implying that this situation has higher average photon numbers than the SA case of the coherent state.
@@ III.3 Impact of multiphoton operations
In §III.1, we noticed that the SA and AS variants of a standard coherent state exhibit nonclassical phase-space features, and it is observed that the sequence in which photon operations are applied significantly affects the outcome, resulting in two distinct quantum states with different phase-space characteristics. We now extended those basic notions to our main quantum states of the present work. First, for our SA scenario, $r$ photons are added to the compass state $\ket{\diamondsuit}$ as given in Eq. (11), followed by the subtraction of $q$ photons. Conversely, in the AS scenario, $q$ photons are first subtracted from the state $\ket{\diamondsuit}$, and then $r$ photons are added.
The mathematical representation of the SA case of the state $\ket{\diamondsuit}$ is
@@
the normalization coefficient.
In the same way, the AS case is obtained as
@@
$N_{\blacksquare}=$ $\sum^{4}_{i,j=1}\sum^{r}_{n=0}(-1)^{n}\Gamma G_{\alpha_{i},\alpha_{j}}\text{e}^{\alpha^{*}_{i}\alpha_{j}}(\alpha^{*}_{i}\alpha_{j})^{q}$
×$\times(\alpha_{j})^{r-n}(\alpha^{*}_{i})^{r-n}$ (27)
represents the normalization factor for this case.
We now compare the original state $\ket{\diamondsuit}$ with the proposed variants by assessing their overlap. This overlap measures distinctness between the proposed states and the original compass state. The overlap between $\ket{\square}$ and $\ket{\diamondsuit}$ is defined as $F_{\square}(c_{0}):=\braket{\square}{\diamondsuit}$. This overlap reads
$F_{\square}(c_{0})=(-\text{i})^{r+q}\left(N_{\square}N_{\diamondsuit}\right)^{\nicefrac{{-1}}{{2}}}\sum^{4}_{i,j=1}F_{\ket{\square_{i}}\bra{\square_{j}}}$ (28)
@@
$F_{\ket{\square_{i}}\bra{\square_{j}}}:=G_{\alpha_{i},\alpha_{j}}\text{e}^{\alpha^{*}_{i}\alpha_{j}}H_{r,q}\left[\text{i}\alpha^{*}_{i},\text{i}\alpha_{j}\right].$ (29)
Similarly, for the AS case, the overlap function $F_{\blacksquare}(c_{0}):=\braket{\blacksquare}{\diamondsuit}$ is
$F_{\blacksquare}(c_{0})=\left(N_{\blacksquare}N_{\diamondsuit}\right)^{\nicefrac{{-1}}{{2}}}\sum^{4}_{i,j=1}F_{\ket{\blacksquare_{i}}\bra{\blacksquare_{j}}},$ (30)
where
$F_{\ket{\blacksquare_{i}}\bra{\blacksquare_{j}}}:=G_{\alpha_{i},\alpha_{j}}\text{e}^{\alpha^{*}_{i}\alpha_{j}}(\alpha^{*}_{i})^{q}(\alpha_{j})^{r}.$ (31)
The corresponding overlaps are shown in Fig. 7 for $r=q=4$, where the red solid line represents the AS case and the blue dashed line represents the overlap SA case. The black dashed line represents the case when $r=q=0$ illustrating the case of the overlap between two same compass states $\ket{\diamondsuit}$. For smaller values of $c_{0}$, the overlaps between the states show a more significant difference, highlighting greater distinctions between the comparable states. However, for larger $c_{0}$ values, the blue dashed and red solid lines converge to the same level as the horizontal black dashed line, indicating that both the SA and AS cases return to the original compass state.
In our work, we stick with the lower $c_{0}$ situation, which essentially belongs to the case of the four-headed kitten state, and then we investigate the corresponding phase-space characteristics of the resultant states obtained by applying a large amount of photon addition (and subtraction) to the indicated four-headed kitten state. Here, note that the AS case of the four-headed kitten state is the eigenstate of the operator $\hat{a}^{4}$, i.e., $\hat{a}^{4}\ket{\blacksquare}=\ket{\diamondsuit}$. This shows that choosing the photon subtraction $q$ as a multiple of 4 ($q=4k$ with $k\in\mathbb{N}^{+}$) corresponds to a photon-added version of the four-headed kitten state. When $p=0$, this reduces to the ordinary case of the four-headed kitten state denoted by $\ket{\diamondsuit}$.
@@ III.4 Unique phase-space features
In the preceding section we introduced the SA and AS cases of our interest, each producing distinct quantum states as observed by their overlap, and now in this section, we particularly employ the Wigner function to investigate the corresponding phase space of these quantum states. To obtain the Wigner function for each case, Eq. (1) is employed, and we denote by $W_{\ket{\square}}(\beta)$ and $W_{\ket{\blacksquare}}(\beta)$ as the corresponding Wigner functions of the SA and AS cases of the kitten state, respectively.
Let us now examine the Wigner distributions for each scenario. For the SA case of our four-headed kitten state introduced in Eq. (24), the Wigner distribution is calculated as
@@
×$\times(\alpha^{*}_{j}-2\beta^{*})^{r-n}.$ (36)
@@
The corresponding Wigner functions are shown in Fig. 8, with Fig. 9 illustrating a close-up of the central phase-space features of each case; Figs. 8(a)-8(c) and 9(a)-9(c) exhibit the SA, while Figs. 8(d)-8(f) and 9(d)-9(f) provide equivalent AS cases of the four-headed kitten state. With specific selections of the parameters $r$ and $q$, Fig. 10 offers a further illustration of the Wigner function of corresponding AS scenarios. It is readily apparent that our SA and AS instances achieve substantially distinct phase-space characteristics as compared to the four-headed kitten state. Interestingly, a central sub-Planck structure is identified in each of the cases outlined. We focus on the significance of the sub-Planck structure in these states and examine this particular phase-space feature in detail.
@@
Photon addition and subtraction appear to have significant effects on the Wigner distribution of the corresponding states. Note that the original four-headed kitten state, represented in Fig. 1(a) with $c_{0}=1$, does not exhibit a sub-Planck structure. This implies that the appearance of the sub-Planck structures in our cases is attributed to the photon operations involved. Specifically, it is observed that as the parameter $r$, which represents the number of added photons, increases, the dimensions of the sub-Planck structure reduce uniformly in both the SA and AS cases and hence can be far smaller than the extension found for the coherent state when $r\gg 1$. This is evident by comparing the scenarios presented in Figs. 9(a) and 9(b) for the SA case, and then Figs. 9(d) and 9(e) are for the corresponding AS cases, where increasing $r$ with constant $q$ clearly depicts this impact. Furthermore, increasing the number of photon subtractions $q$ increases the size of the sub-Planck structure in the SA case, as observed by comparing cases depicted in Figs. 8(b) and 8(c), while it has no effect in the AS case, as shown in Figs. 8(e) and 8(f). Note that all the examples of the AS described in Figs. 8 and 9 correspond to the merely photon-added case of the compass state, as for these situations $q$ is a multiple of 4, where the corresponding four-headed kitten state is the eigenstate of the operator $\hat{a}^{4}$.
Figure 10: Close-up of the center phase-space structures of the AS case of the four-headed kitten state with multiple values of the parameters $r$ and $q$: (a) $r=1$ and $q=1$, (b) $r=5$ and $q=5$, (c) $r=9$ and $q=9$, (d) $r=16$ and $q=10$, (e) $r=22$ and $q=10$, (f) $r=22$ and $q=18$. For all cases, $c_{0}=1$.
@@
We find that photon addition generally reduces the size of sub-Planck structures in both SA and AS scenarios, while photon subtraction increases their size in the SA case but has no effect on their size in the AS case. To further illustrate this for the AS case, we examine situations where the number of photon subtractions is not a multiple of 4, as shown in Fig. 10. Increasing the photon addition to the state as previously noted also reduces the size of the sub-Planck structures for this specific illustration [Figs. 10(d) and 10(e)] while as expected, varying the photon subtraction in these cases has no impact on the size of the sub-Planck structures [Figs. 10(e) and 10(f)], but in some specific instances, photon subtraction slightly enhances the isotropy of the sub-Planck structures. For example, an enhancement in the isotropy of the central sub-Planck structures of these cases is clearly evident in Figs. 10(a) to 10(c), as demonstrated by the circular structure around the origin, which is further analyzed in the Appendix A. This indicates that, for these scenarios, the sub-Planck structure is uniformly constrained in all directions of the phase space, hence representing an isotropic version of the sub-Planck structures (see the Appendix A).